• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Deur

    Log entry time 00:13:47 on October 5,2002

    Entry number 87338

    keyword=Shift Summary

    5:00 Q1 left arm fixed
    5:20: Start sweeping the Hall in order to go in beam permit
    5:50. Set the left HRS to P0=2.011 GeV/c (cycled Q2 & Q3). Meanwhile we moved left HRS to 38.35 degree
    6:24: Beam tuning is done. We moved the target from empty target to the BeO viewer foil. We requested MCC for 4 uA and raster off.
    6:46. Beam here: 4.4 uA unrastered. But beam is not stable enough for an harp scan. MCC said that it is because the moller quads were not setup correctly. They also mention a problem with the fast feadback (now turned off)
    The Moller quads are setting now the correct optics.
    7:20 Beam is usable. Bodo starts the harp scan
    Trouble with harp scan (one of the harp is not working).
    Asked back for the raster (2*2 mm)
    Took runs 1129-1132 with roc14 (needed for spot++) while Bodo is scanning the harp.
    20:11 Put the C12 optic target in.
    20:15 Ask for 30 uA of rastered beam (2*2 mm)
    20:22. while waiting for beam, MCC called saying they need to cycle a magnet.
    20:44 Set the left HRS to 2.775 GeV/c (cycled q2 and q3)
    mcc said they cannot send us high current ( I>5 uA) without tripping the ion chamber 1H03B right. (Beam dump ion chamber).
    Q1 right tripped. Need an access to fix it.
    The cryotarget warmed up a little bit (19.2 degree) with the power heater load going to 0. -> Opened the JT valves from 27 to 31. The ESR flow increased a bit and then the cryotarget warmed up a bit. I don't know if it is related. The temperature of the coolant was stable.
    Still problem with getting the beam without tripping the ion chamber.
    22:23 MCC is done with the Hall A tunning. We still don't have beam.
    23:50 Moved the target to empty position at MCC request.